OpenStudy (anonymous):

if 13/4 equal 50 miles how many miles equal 7 inches

OpenStudy (anonymous):

I would change the improper fraction to a decimal. So 13/4 = 3.25 Next I would write the problem as a proportion. So it would look like this: \[\frac{ 3.25 inches }{ 50miles}=\frac{ 7 inches }{ ? miles}\] Cross multiply 7 times 50, which gives us 350 And divide by the other number.(3.25) 350/3.25=107.69 So the ? = 107.69 If we round to the nearest mile, the answer is 108 miles.